### v1

@@ 6.4 Hyperparameters
-0.1222This is a log-probability. -0.02222This is a log-probability.
$V_{\max}$ Probability
@@ 7 Results and Discusion
The results of the Threshold utility function for the Quay Wall Asset environment show an even lower cost with MO-DCMAC (0.931 ± 0.15) compared to CBI-CBA (1.479 ± 0.236); however, the probability of collapse of MO-DCMAC (0.028 ± 0.033) is significantly higher than of CBI-CBA (0.046 ± 0.052). This difference is because MO-DCMAC learned it should not do certain maintenance actions if it decreases the probability of collapse beyond the threshold. Nevertheless, we see in Table 2 that the probability of collapse is, on average, always under the threshold of 0.1. This is true for all the environments, which indicates that during the training process, MO-DCMAC wants to learn a policy that can also handle unforeseen failures that can suddenly increase the probability of collapse. Still, it does not consistently achieve a probability of collapse below the threshold, with the utility score always being higher than the cost. This effect is especially noticeable in the Quay Wall Asset environment with MO-DCMAC (1.133 ± 1.186) because the standard deviation is almost equal to the mean utility score, while this is not the case with the other policies. Therefore, with MO-DCMAC, the probability of being above the threshold is higher than with other policies. A solution might be to increase the penalty if the probability of collapse is above the threshold.
@@
Figure 6: The figure shows the effect of using a different $\gamma$ with the Threshold utility function. The blue bar is the non-discounted value, and the orange bar is the discounted value, which is the value if we apply the corresponding $\gamma$. For each of the plots, a lower score is always strictly better.
In Section 6.2.1, we stated how we wanted to show the effect on the Threshold utility with different discount factors $\gamma$, namely $\gamma=0.9$, $\gamma=0.975$, $\gamma=0.99$, $\gamma=0.995$, and $\gamma=1$. We argued that a low discount factor might result in a sub-optimal policy due to the difference between the discounted and non-discounted return can have have on the utility score.
@@ Appendix A Environments
$\forall k\in K\wedge\forall i\in\{1,...,\tau_{\max}-1\}$ (27)
@@ A.2 Quay Wall Asset
$\operatorname{\mathcal{G}}_{1}^{\text{poles}}$ =$=\{\operatorname{c}_{1},\operatorname{c}_{2},\operatorname{c}_{3}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{poles}}$ =$=\{\operatorname{c}_{4},\operatorname{c}_{5},\operatorname{c}_{6}\}$
$\operatorname{\mathcal{G}}_{3}^{\text{poles}}$ =$=\{\operatorname{c}_{7},\operatorname{c}_{8},\operatorname{c}_{9}\}.$ (39)
@@
$\operatorname{\mathcal{G}}_{1}^{\text{kesps}}$ =$=\{\operatorname{c}_{10},\operatorname{c}_{11}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{kesps}}$ =$=\{\operatorname{c}_{11},\operatorname{c}_{12}\}.$ (40)
@@
$\operatorname{\mathcal{G}}_{1}^{\text{floor}}=\{\operatorname{c}_{13}\}.$ (41)
@@ A.3 Larger Quay Wall Asset
The last environment, Larger Quay Wall, is similar to the previous environment, Quay Wall, with the major difference being that the number of components is doubled to $N=26$. Therefore, we have the same component groups of wooden poles, kesps, and floors. The wooden poles are $\{\operatorname{c}_{1},...,\operatorname{c}_{18}\}$, with the degradation matrices found in Eqs. 33 and 34, the wooden kesps are $\{\operatorname{c}_{19},...,\operatorname{c}_{24}\}$ with the degradation matrices found in Eqs. 35 and A.2, and the wooden floors are $\{\operatorname{c}_{25},\operatorname{c}_{26}\}$ with the degradation matrices found in Eqs. 35 and A.2. The cost of each maintenance action is found in Table 7.
@@
$\operatorname{\mathcal{G}}_{1}^{\text{poles}}$ =$=\{\operatorname{c}_{1},\operatorname{c}_{2},\operatorname{c}_{3}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{poles}}$ =$=\{\operatorname{c}_{4},\operatorname{c}_{5},\operatorname{c}_{6}\}$
$\operatorname{\mathcal{G}}_{3}^{\text{poles}}$ =$=\{\operatorname{c}_{7},\operatorname{c}_{8},\operatorname{c}_{9}\}$
$\operatorname{\mathcal{G}}_{4}^{\text{poles}}$ =$=\{\operatorname{c}_{10},\operatorname{c}_{11},\operatorname{c}_{12}\}$
$\operatorname{\mathcal{G}}_{5}^{\text{poles}}$ =$=\{\operatorname{c}_{13},\operatorname{c}_{14},\operatorname{c}_{15}\}$
$\operatorname{\mathcal{G}}_{6}^{\text{poles}}$ =$=\{\operatorname{c}_{16},\operatorname{c}_{17},\operatorname{c}_{18}\}.$ (43)
@@
$\operatorname{\mathcal{G}}_{4}^{\text{kesps}}$ =$=\{\operatorname{c}_{22},\operatorname{c}_{23}\}$
$\operatorname{\mathcal{G}}_{5}^{\text{kesps}}$ =$=\{\operatorname{c}_{23},\operatorname{c}_{24}\}.$ (44)
@@
$\operatorname{\mathcal{G}}_{1}^{\text{floor}}$ =$=\{\operatorname{c}_{25}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{floor}}$ =$=\{\operatorname{c}_{26}\}.$ (45)

### v2

@@ 6.4 Hyperparameters
-0.1 -0.02
$V_{\max}$ Log Probability
@@ 7 Results and Discusion
The results of the Threshold utility function for the Quay Wall Asset environment show an even lower cost with MO-DCMAC (0.931 ± 0.15) compared to CBI-CBA (1.479 ± 0.236); however, the probability of collapse of MO-DCMAC (0.028 ± 0.033) is significantly higher than of CBI-CBA (0.01 ± 0.018). This difference is because MO-DCMAC learned it should not do certain maintenance actions if it decreases the probability of collapse beyond the threshold. Nevertheless, we see in Table 2 that the probability of collapse is, on average, always under the threshold of 0.1. This is true for all the environments, which indicates that during the training process, MO-DCMAC wants to learn a policy that can also handle unforeseen failures that can suddenly increase the probability of collapse. Still, it does not consistently achieve a probability of collapse below the threshold, with the utility score always being higher than the cost. This effect is especially noticeable in the Quay Wall Asset environment with MO-DCMAC (1.133 ± 1.186) because the standard deviation is almost equal to the mean utility score, while this is not the case with the other policies. Therefore, with MO-DCMAC, the probability of being above the threshold is higher than with other policies. A solution might be to increase the penalty if the probability of collapse is above the threshold.
@@
Figure 6: The figure shows the effect of using a different $\gamma$ with the Threshold utility function with the Simple Homogenous Asset environment. The blue bar is the non-discounted value, and the orange bar is the discounted value, which is the value if we apply the corresponding $\gamma$. For each of the plots, a lower score is always strictly better.
In Section 6.2.1, we stated how we wanted to show the effect on the Threshold utility with different discount factors. We argued that a low discount factor might result in a sub-optimal policy due to the difference between the discounted and non-discounted returns, which can have an effect on the utility score. For these experiments, we used the Simple Homogenous Asset environment with the following discount factors: $\gamma=0.9$, $\gamma=0.975$, $\gamma=0.99$, $\gamma=0.995$, and $\gamma=1$
@@ Appendix A Environments
$\forall k\in K\wedge\forall i\in\{0,...,\tau_{\max}-1\}$ (27)
@@ A.2 Quay Wall Asset
$\operatorname{\mathcal{G}}_{0}^{\text{poles}}$ =$=\{\operatorname{c}_{0},\operatorname{c}_{1},\operatorname{c}_{2}\}$
$\operatorname{\mathcal{G}}_{1}^{\text{poles}}$ =$=\{\operatorname{c}_{3},\operatorname{c}_{4},\operatorname{c}_{5}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{poles}}$ =$=\{\operatorname{c}_{6},\operatorname{c}_{7},\operatorname{c}_{8}\}.$ (39)
@@
$\operatorname{\mathcal{G}}_{0}^{\text{kesps}}$ =$=\{\operatorname{c}_{9},\operatorname{c}_{10}\}$
$\operatorname{\mathcal{G}}_{1}^{\text{kesps}}$ =$=\{\operatorname{c}_{10},\operatorname{c}_{11}\}.$ (40)
@@
$\operatorname{\mathcal{G}}_{0}^{\text{floor}}=\{\operatorname{c}_{12}\}.$ (41)
@@ A.3 Larger Quay Wall Asset
The last environment, Larger Quay Wall, is similar to the previous environment, Quay Wall, with the major difference being that the number of components is doubled to $N=26$. Therefore, we have the same component groups of wooden poles, kesps, and floors. The wooden poles are $\{\operatorname{c}_{1},...,\operatorname{c}_{18}\}$, with the degradation matrices found in Equations 33 and 34, the wooden kesps are $\{\operatorname{c}_{19},...,\operatorname{c}_{24}\}$ with the degradation matrices found in Equations 35 and A.2, and the wooden floors are $\{\operatorname{c}_{25},\operatorname{c}_{26}\}$ with the degradation matrices found in Equations 35 and A.2. The cost of each maintenance action is found in Table 7.
@@
$\operatorname{\mathcal{G}}_{0}^{\text{poles}}$ =$=\{\operatorname{c}_{0},\operatorname{c}_{1},\operatorname{c}_{2}\}$
$\operatorname{\mathcal{G}}_{1}^{\text{poles}}$ =$=\{\operatorname{c}_{3},\operatorname{c}_{4},\operatorname{c}_{5}\}$
$\operatorname{\mathcal{G}}_{2}^{\text{poles}}$ =$=\{\operatorname{c}_{6},\operatorname{c}_{7},\operatorname{c}_{8}\}$
$\operatorname{\mathcal{G}}_{3}^{\text{poles}}$ =$=\{\operatorname{c}_{9},\operatorname{c}_{10},\operatorname{c}_{11}\}$
$\operatorname{\mathcal{G}}_{4}^{\text{poles}}$ =$=\{\operatorname{c}_{12},\operatorname{c}_{13},\operatorname{c}_{14}\}$
$\operatorname{\mathcal{G}}_{5}^{\text{poles}}$ =$=\{\operatorname{c}_{15},\operatorname{c}_{16},\operatorname{c}_{17}\}.$ (43)
@@
$\operatorname{\mathcal{G}}_{0}^{\text{kesps}}$ =$=\{\operatorname{c}_{18},\operatorname{c}_{19}\}$
@@
$\operatorname{\mathcal{G}}_{4}^{\text{kesps}}$ =$=\{\operatorname{c}_{22},\operatorname{c}_{23}\}.$ (44)
@@
$\operatorname{\mathcal{G}}_{0}^{\text{floor}}$ =$=\{\operatorname{c}_{24}\}$
$\operatorname{\mathcal{G}}_{1}^{\text{floor}}$ =$=\{\operatorname{c}_{25}\}.$ (45)